4.5
Solve the equation x^3 -13x^2 +47x-35=0 given that 1 is a zero of f(x)=x^3 -13x^2 +47x-35
The solution set is { }

I got 1,5,7 is this right?

Answer by Alan3354(69443)

Solve the equation x^3 -13x^2 +47x-35=0 given that 1 is a zero of f(x)=x^3 -13x^2 +47x-35
=======================
I got 1,5,7 is this right?
------------
It's right.
You can check it.
If (x-1)*(x-5)*(x-7) = x^3 -13x^2 +47x-35 it's right.
---
f(x) = x^3 -13x^2 +47x-35
If f(1) = 0 it's a zero.
Same for f(5) and f(7).